OpenStudy (anonymous):

Solve the following system using the substitution method. x + 6y = -5 2x - 9y = -17

OpenStudy (anonymous):

Anyone know how to do this?

OpenStudy (anonymous):

1) x=-6y-5 2) 2x-9y=-17 substitute (1) in (2) 2(-6y-5)-9y=-17 -12y-10-9y=-17 -21y=10-17 -21y=-7 3)\[y=\frac{ 1 }{ 3}\] substitute (3) in (1): \[x=-6(\frac{ 1 }{ 3 })-5=-7\]

OpenStudy (anonymous):

What is the best first step to solve the system using elimination? x - 5y = 4 3x + 7y = -17

OpenStudy (anonymous):

multiply fist equation by 3 and then substract them